OpenStudy (anonymous):

The rate traveled from Amarillo to Austin by a bus averages 60 miles per hour. The bus arrived in Austin after eight hours of travel. An automobile averages 48 miles per hour. Using the inverse variation relationship, show what the time would be for the automobile to complete the trip.

OpenStudy (cwrw238):

distance is the same so speed = distance / time 60 = d /t = d /8 d = 60*8 miles car travels at 48 mph 48 = d / t = 60*8 / t t = 60*8 / 48 = 10 hours